OpenStudy (anonymous):

Factor the algebraic expression below in terms of a single trigonometric function. cos x - sin ^2x - 1

OpenStudy (anonymous):

\[\sin^{2}x = 1- \cos^{x}\] so \[cosx - (1-\cos^{2}x) -1 = \cos^{2}x + cosx -2\]

OpenStudy (anonymous):

let x = cosx now you need to factor \[x^{2} + x -2 = 0\]

OpenStudy (anonymous):

Ok thank you